exercise:
write a formula for the nth term of each arithmetic sequence:
1. -4,-2,0,2,4,...
2. 5,1,-3,-7,-11,..
how can i solve these?
can you help me?
and could you please explain it to me how it come so that i can understand it better...
thank you very much...
your responce will highly appreciated..

Answer by mathslover(157)

nth term of an arithmetic sequence is give by
t(n) = a + (n-1)d where a is the first team of the sequence and d the common difference
for the first series
a=-4
d= -2 - (-4) difference between any 2 successive terms
= 2
t(n) = -4 + (n-1)2
= -4 + 2n -2
= 2n - 6
for the second series
a=5
and d = 1-5 = -4
the nth term
t(n) = 5 + (n-1)* -4
= -4n +9